# Liquid: Language Models are Scalable Multi-modal Generators

### v2

@@ Abstract
We present Liquid, an auto-regressive generation paradigm that seamlessly integrates visual comprehension and generation by tokenizing images into discrete codes and learning these code embeddings alongside text tokens within a shared feature space for both vision and language. Unlike previous multimodal large language model (MLLM), Liquid achieves this integration using a single large language model (LLM), eliminating the need for external pretrained visual embeddings such as CLIP. For the first time, Liquid uncovers a scaling law that performance drop unavoidably brought by the unified training of visual and language tasks diminishes as the model size increases. Furthermore, the unified token space enables visual generation and comprehension tasks to mutually enhance each other, effectively removing the typical interference seen in earlier models. We show that existing LLMs can serve as strong foundations for Liquid, saving 100× in training costs while outperforming Chameleon in multimodal capabilities and maintaining language performance comparable to mainstream LLMs like LLAMA2. Liquid also outperforms models like SD v2.1 and SD-XL (FID of 5.47 on MJHQ-30K), excelling in both vision-language and text-only tasks. This work demonstrates that LLMs such as Qwen2.5 and GEMMA2 are powerful multimodal generators, offering a scalable solution for enhancing both vision-language understanding and generation. The code and models will be released at https://github.com/FoundationVision/Liquid.
@@ 1 Introduction
The recent advancement of Large Language Models (LLMs) [4, 64, 8, 48, 65, 61] has sparked a trend of extending the foundational capabilities of LLMs to the visual domain, giving rise to Multi-modal Large Language Models (MLLMs) [39, 77, 37, 42, 5, 38]. In the realm of visual understanding, MLLMs such as LLaVA [39] typically adopt the pretrained CLIP [46] model as the visual tokenizer, followed by a two-stage training process to align the vision-and-language feature space. While for text-guided visual generation, most MLLMs [18, 19, 25, 59, 58] rely on an external diffusion model to generate images. Despite these methods have achieved remarkable multi-modal understanding and generation performances, the use of external visual modules introduces additional architecture complexity to the system, and potentially poses a bottleneck when scaling up the LLMs.
In light of the aforementioned issues, an emerging line of research [15, 50, 12, 72, 63, 57] attempts to employ the VQVAE [66, 15] model as a universal visual tokenizer for MLLMs. Analogous to the role played by the BPE [54] tokenizer in LLMs, VQVAE establishes a bi-directional mapping between raw pixels and discrete codes. This enables the MLLMs to learn visual code embeddings jointly with text tokens, rather than constrained by the feature space of a pretrained visual encoders like CLIP or pretrained visual generators like diffusion models. Moreover, the discrete nature of visual tokens allows uniform modeling of visual and text tokens with the same next-token prediction loss, which seamlessly integrates both modalities. LWM [40] and Chameleon [60] are the pioneers in exploring this approach. However, these methods involve extensive training from scratch, which makes it computationally expensive to disseminate exploration in this form. Follow-up works further introduces bi-directional attention [76], discrete diffusion modeling [71], and residual codes [70] to handle visual tokens. EMU3 [67] has verified that the next prediction paradigm can achieve SOTA results in both understanding and generation tasks by finetuning separate two models, but there is a lack of exploration of the effects of unifying these two tasks within one model. In this paper, we explore the potential of solely employing LLMs as multimodal generators, and conduct a comprehensive empirical study to investigate a series of properties including the scaling laws of visual generation, the impact on original language capabilities, and the interrelation between visual understanding and generation tasks.
@@
Our experiments reveal several insightful properties: 1) Directly employing LLMs for visual generation exhibits clear scaling laws in both validation loss and image consistency metrics, consistent with those seen in LLMs, regardless of whether the models retain their language capabilities or focus solely on visual generation tasks. 2) When trained with multimodal data such as image-text pairs and pure language data, the language capabilities of the models are lower compared to those trained with only language data. However, this tradeoff diminishes as the model size increases, indicating that larger models have sufficient capacity to handle both tasks seamlessly. This, combined with the scaling law, demonstrates the strong potential of LLMs as unified multimodal generator. 3) The visual understanding and visual generation tasks can mutually benefit each other. We find that increasing the data for either visual generation or visual understanding tasks improves the performance of the other, demonstrating the advantages and natural rationale of this paradigm in joint optimization.
@@
  • Comprehensive experiments about scaling laws of unified multi-modal models, revealing that the trade-off between language and visual tasks diminishes as scale of model increases.
@@
2 Preliminaries
Image Tokenizer. We use the same tokenizer as Chameleon [60], which is a VQGAN [15]. It encodes a 512 × 512 image into 1024 discrete tokens in a codebook of size 8192. We appended these discrete image tokens to the text codebook produced by the BPE [54] tokenizer to expand the vocabulary size of the LLMs, thereby extending its language space to a multi-modal space encompassing both visual and linguistic elements.
Architecture. As illustrated in Fig. 1, Liquid can be constructed based on any existing LLMs. In this paper, we employ the GEMMA-7B [61] as base model of Liquid to validate its multimodal understanding, image generation capabilities, and performance on text-only tasks after augmenting it with the ability to understand and generate images. Then we conduct a comprehensive empirical study on the scaling behavior at scales of 0.5B, 1B, 2B, 7B, 9B, and 32B models from LLAMA-3 [14], GEMMA-2 [62], and Qwen2.5 [24]. We refrain from altering any structures within the LLMs to facilitate continued training directly from pre-trained weights. The only modification is the addition of 8192 new learnable embeddings for discrete image tokens. The LLMs maintain their original next-token prediction training objective with the standard cross-entropy loss.
Data Preparation. To maintain the language abilities of pre-trained LLMs, we sample text-only data from public datasets for the pre-training stage. We sample 15M text data from DCLM [28], 12M from SlimPajama [56], and 3M code data from Starcoderdata [31], totaling 30M text instances for approximately 60 billion text tokens. For image-text pairs, we use JourneyDB [44] and internal MidJourney-style synthetic data to compile 30M high-quality image data, for a total of 30 billion image tokens. All the data are used to form a hybrid multi-modal data for continue pre-training, resulting in quick acquisition of decent image generation capabilities while retaining language abilities. Furthermore, we reversed 20% of text-to-image data to train LLMs in a captioning task to enhance their visual understanding abilities.
Training Procedure. We use a total of 60M data to continue pretraining our models. For all image-text data, we define the input format for multi-modal training data as:
@@ 2 Preliminaries
Here, [bos] and [eos] are the begin-of-sequence and end-of-sequence tokens defined in the original text tokenizer. We incorporate two additional spatial tokens, [boi] and [eoi] to signify the start and end of image tokens. In the scaling experiments, for each model size, we independently use 30M text-only data, 30M text-to-image data, and a mixed dataset of 60M data to train three different versions of the models. We then evaluate their performance on a series of tasks.
@@ 3 Experiments
3.1 How Well Do LLMs Perform in Visual Generation?
In this section, we comprehensively evaluate the visual generation capabilities of Liquid-7B model that have been continuously pretrained with a combined dataset of 60M mixed multimodal data. Subsequently, we evaluate mixed pre-trained models across six different sizes, ranging from 0.5B to 32B, the objective is to explore the scaling behavior of LLMs in visual generation tasks, while simultaneously preserving their linguistic capabilities.
@@ 3.1 How Well Do LLMs Perform in Visual Generation?
3.1.1 Quantitative Results on Visual Generation Benchmarks
Table 2: Comparison with other visual generation methods on MJHQ-30K evaluation benchmark. The FID of Liquid is lower than that of all the auto-regressive models and even outperforms most diffusion models. This indicates that the images generated by Liquid have superior aesthetic quality.
@@ 3.1.1 Quantitative Results on Visual Generation Benchmarks
For the image generation tasks, we evaluate the model on two benchmarks, GenAI-Bench [36] and MJHQ-30K [27]. GenAI-Bench is an challenging image-to-text generation benchmark designed to evaluate the capabilities of visual generation models. It evaluates both basic skills, such as understanding attributes, scenes, and relations in text inputs, and advanced skills, including abilities like counting, differentiation, comparison, and understanding logical relationships from text input, thereby highlighting the graded progression in complexity. GenAI-Bench employs VQAScore, which leverages a visual-question-answering (VQA) model. This enables more precise evaluation of how well the generated image aligns with the text prompt, critically assessing the capability to parse scenes, objects, attributes, relationships, and engage in higher-order reasoning such as comparison and logic. The latter calculates the Frechet Inception Distance (FID) [22] score between the generated images and 30K high-quality images to assess the quality of the generated images.
@@
As shown in Tab. 3.1, compare with other auto-regressive based methods, Liquid achieves a better overall score under both basic prompts and advanced prompts. This suggests that the images generated by Liquid align better semantically with the input text prompts. Notably, Liquid also outperforms some well-established diffusion models like SD v2.1 [51] and SD-XL[45] for both basic and advanced prompts. Compared to these diffusion models, Liquid utilizes significantly fewer image data, which indicating that learning based on LLMs can assist the model in understanding the semantic association between the generated content and prompts, while also offering higher training efficiency. Moreover, it demonstrates that LLMs have strong potential for generating complex visual content.
@@
In Tab. 2, we report FID on MJHQ-30K to compare the images quality generated by Liquid with other models. It is observable that Liquid not only has a lower FID than all other auto-regressive methods but also surpasses most well-known diffusion models except Playground v2.5 [27], achieving a very low FID of 5.47. It indicates that LLMs are also capable of generating high-quality images, as shown in Fig. 2, providing proof that the upper limit of LLMs in terms of image aesthetic quality is not inferior to diffusion models. Further more, due to the capability of LLM to generate dynamically-length content in the form of next-token prediction, the convenience can be applied to visual generation. We find that by appending instructions about the resolution to the input text prompt, such as "length is: width is:", the model can quickly learn to generate the corresponding code according to the specified number of rows and columns. Fig. 2 demonstrates the generation results at various resolutions, showcasing the flexibility of Liquid.
Figure 3: Scaling behavior of LLMs in multimodal generation across different sizes. We explore the visual generation performance of LLMs ranging from 0.5B to 32B in size after undergoing mixed training with language data and text-to-image data. The validation loss smoothly decreases as both the model size and training iterations increase, while token accuracy and VQA Score consistently rise. The VQA Score consistently increases as the validation loss decreases, indicating a strong correlation between them.Uder the same training FLOPs, smaller models reach lower validation losses more quickly, but larger models ultimately achieve higher evaluation metrics.
3.1.2 Scaling Results on Visual Generation
We explore the visual generation performance of LLMs ranging from 0.5B to 32B in size after mixed training with language data and text-to-image data. As shown in Fig. 3, with the increase in model size and training iterations, the validation loss smoothly decreases, while token accuracy and VQA Score consistently increase. Moreover, the VQA score continues to rise as the validation loss decreases, indicating a strong correlation between validation loss and holistic image evaluation metrics. Under the same training FLOPs, smaller models achieve lower validation losses and higher VQA Scores more quickly, but larger models ultimately reach higher evaluation metrics. This may be due to smaller models being able to iterate through more training steps rapidly, allowing for swift changes in model capabilities. Although smaller models can quickly acquire visual generation abilities, their upper limits are lower, making it difficult to achieve high-quality visual generation results. As shown in Fig. 4, larger models eventually obtain more robust visual generation outcomes.
3.2 Does Visual Generation Compromise Language Ability?
In this section, we evaluate the impact on the inherent linguistic capabilities of LLMs after they acquire visual generation abilities. First, we compare the continuously pretrained GEMMA-7B base model with mainstream language models to determine if it has retained robust linguistic capabilities. Subsequently, we evaluate LLMs of different sizes that have been trained either exclusively with 30M language data or with a mix of 60M multimodal data. Since both training processes use the same language data, this allows for a fair comparison of how different sizes of LLMs are affected in terms of their original linguistic capabilities after acquiring visual generation abilities.
3.2.1 Comparison with Mainstream LLMs
To validate whether acquiring image understanding and generation capabilities has any impact on the original language abilities of the LLMs, we report overall zero-shot performance across a suite of popular benchmarks that measure commonsense reasoning and reading comprehension capabilities: HellaSwag [74], WinoGrande [52] , ARC-Easy [10], ARC-Challenge [10], OpenBookQA [43], PIQA [3], SIQA [53], and BoolQ [9]. We also perform an evaluation of the 5-shot results on MMLU [21], a comprehensive benchmark that measures world/in-domain knowledge and problem-solving skills across 57 subjects.
We evaluate the general text-only capabilities of our mixed multimodal pretrained model against other state-of-the-art large language models and multi-modal language models, following the evaluation protocol of  [17]. As shown in Tab. 3, Liquid outperforms the well-established language model LLAMA2 [65] and the mix-pretrained multi-modal language model Chameleon [60] in most tasks, exhibiting undegraded linguistic capabilities. Compared with Chameleon [60], which is mixed pretrained with an extremely large scale of data, Liquid performs training from existing LLMs that already possess decent language capabilities, maintaining these capabilities without degradation. This result validates the efficiency of our training framework and demonstrates that with this framework, we can extend the visual generation and understanding capabilities to LLMs of any structure and size.

### v4

# Liquid: Language Models are Scalable and Unified Multi-modal Generators

@@ Abstract
We present Liquid, an auto-regressive generation paradigm that seamlessly integrates visual comprehension and generation by tokenizing images into discrete codes and learning these code embeddings alongside text tokens within a shared feature space for both vision and language. Unlike previous multimodal large language model (MLLM), Liquid achieves this integration using a single large language model (LLM), eliminating the need for external pretrained visual embeddings such as CLIP. For the first time, Liquid uncovers a scaling law that performance drop unavoidably brought by the unified training of visual and language tasks diminishes as the model size increases. Furthermore, the unified token space enables visual generation and comprehension tasks to mutually enhance each other, effectively removing the typical interference seen in earlier models. We show that existing LLMs can serve as strong foundations for Liquid, saving 100× in training costs while outperforming Chameleon in multimodal capabilities and maintaining language performance comparable to mainstream LLMs like LLAMA2. Liquid also outperforms models like SD v2.1 and SD-XL (FID of 5.47 on MJHQ-30K), excelling in both vision-language and text-only tasks. This work demonstrates that LLMs such as Qwen2.5 and GEMMA2 are powerful multimodal generators, offering a scalable solution for enhancing both vision-language understanding and generation. The code and models are released at https://github.com/FoundationVision/Liquid.
@@ 1 Introduction
Figure 1: Compared to existing unified MLLMs, Liquid eliminates the need for additional model composition, semantic pre-training, or separate output heads. It directly performs next-token prediction and training within a unified vision-language space, requiring only an extension of vocabulary and low-cost continue training. This design enables seamless adaptation to any existing LLMs.
The recent advancement of Large Language Models (LLMs) [4, 69, 8, 52, 70, 65] has sparked a trend of extending the foundational capabilities of LLMs to the visual domain, giving rise to Multi-modal Large Language Models (MLLMs) [39, 83, 37, 42, 5, 38]. In the realm of visual understanding, MLLMs such as LLaVA [39] typically adopt the pretrained CLIP [50] model as the visual tokenizer, followed by a two-stage training process to align the vision-and-language feature space. While for text-guided visual generation, most MLLMs [17, 18, 24, 63, 62, 68] rely on an external diffusion model to generate images as shown in Fig. 1(a). Despite these methods have achieved remarkable multi-modal understanding and generation performances, the use of external visual modules introduces additional architecture complexity to the system, and potentially poses a bottleneck when scaling up the LLMs.
In light of the aforementioned issues, an emerging line of research [15, 54, 12, 78, 67, 61] attempts to employ the VQVAE [71, 15] model as a universal visual tokenizer for MLLMs. Analogous to the role played by the BPE [58] tokenizer in LLMs, VQVAE establishes a bi-directional mapping between raw pixels and discrete codes. This enables the MLLMs to learn visual code embeddings jointly with text tokens, rather than constrained by the feature space of a pretrained visual encoders like CLIP or pretrained visual generators like diffusion models. Moreover, the discrete nature of visual tokens allows uniform modeling of visual and text tokens with the same next-token prediction loss, which seamlessly integrates both modalities. LWM [40] and Chameleon [64] are the pioneers in exploring this approach. However, these methods involve extensive training from scratch, which makes it computationally expensive to disseminate exploration in this form. Follow-up works further introduce diffusion modeling [82, 76] but leads to inconsistent training objectives between visual and text generation as show in Fig. 1(b). (c) and (d) require additional semantic pretraining for image tokenizers [74, 75, 49], while (e) necessitates additional preprocessing and postprocessing steps for visual tokens [27]. EMU3 [72] has verified that the next prediction paradigm can achieve SOTA results in both understanding and generation tasks by finetuning separate two models, but there is a lack of exploration of the effects of unifying these two tasks within one model. In this paper, we explore the potential of solely employing LLMs as multimodal generators, and conduct a comprehensive empirical study to investigate a series of properties including the scaling laws of visual generation, the impact on original language capabilities, and the interrelation between visual understanding and generation tasks.
@@
Our experiments reveal several insightful properties: 1) Directly employing LLMs for visual generation exhibits clear scaling laws in both validation loss and image consistency metrics, consistent with those seen in LLMs, regardless of whether the models retain their language capabilities or focus solely on visual generation tasks. 2) When trained with multimodal data such as image-text pairs and pure language data, the language capabilities of the models are lower compared to those trained with only language data. However, this tradeoff diminishes as the model size increases, indicating that larger models have sufficient capacity to handle both tasks seamlessly, demonstrating the strong potential of LLMs as unified multimodal generator. 3) The visual understanding and visual generation tasks can mutually benefit each other. We find that increasing the data for either visual generation or visual understanding tasks improves the performance of the other, demonstrating the advantages and natural rationale of this paradigm in joint optimization.
@@
  • Comprehensive experiments about scaling laws of unified autoregressive multi-modal models, revealing that the trade-off between language and visual tasks diminishes as scale of model increases.
@@
2 Preliminaries
Image Tokenizer. We use a VQGAN [15] from Chameleon [64] as image tokenizer. It encodes a 512 × 512 image into 32 × 32 discrete tokens in a codebook of size 8192, which are flattened into a 1024 image token for LLM. In LLMs such as Llama [69] and Gemma [65], the raw text input will also be tokenized into numbers (discrete IDs) representing the positions of the words in the vocabulary by the BPE [58] tokenizer. For example, "a small dog" is tokenized into [235250, 2301, 5929] in Gemma, the vocabulary size is 256,000. Therefore, we only need to extend the entire vocabulary size to $256,000+8192$, where the former for text and the latter for image. In this way, both vision and language are represented as sequences of discrete IDs, whether for input or output. After that, IDs index the input embeddings through an embedding-layer, thus completing the unification of the embedding space. Correspondingly, we extend the original LM head by 8192 dimensions to enable the model to predict both text and image tokens within the same embedding space.
@@ 2 Preliminaries
Architecture. As illustrated in Fig. 2, Liquid can be constructed based on any existing LLMs. In this paper, we employ the GEMMA-7B [65] as base model of Liquid to validate its multimodal understanding, image generation capabilities, and performance on text-only tasks after augmenting it with the ability to understand and generate images. To further investigate whether a trade-off exists when accommodating visual generation tasks and text generation tasks within the same LLM space, as well as to explore the scaling performance of different sized models, we conduct same training on Llama-3 1B  [14], GEMMA-2 series [66] at scales of 2B, 9B, and Qwen2.5 series [23] at scales of 0.5B, 7B, 32B, thereby observing their distinct performances. We refrain from altering any structures within the LLMs to facilitate continued training directly from pre-trained weights. The only modification is the addition of 8192 new learnable embeddings for discrete image tokens. Correspondingly, we extend the original LM head by 8192 dimensions to enable the model to predict both text and image tokens within the same embedding space. The LLMs maintain their original next-token prediction training objective with cross-entropy loss. , with the only variation being that their vocabulary space has been extended from purely language-space to a combined space for both visual and linguistic modalities.
Data Preparation. To maintain the language abilities of the pre-trained LLMs, we sample text-only data from public datasets for the pre-training stage. We sample 15M text data from DCLM [28], 12M from SlimPajama [60], and 3M code data from Starcoderdata [31], totaling 30M text instances for approximately 60 billion text tokens. For image-text pairs, we use JourneyDB [46] and internal MidJourney-style synthetic data to compile 30M high-quality image data, for a total of 30 billion image tokens. All the data are used to form a hybrid multi-modal data for continue pre-training, resulting in quick acquisition of image generation capabilities while retaining language abilities.
Training Procedure. We use a total of 60M data to continue pretraining our models. For all image-text pair data, we define the input format for multi-modal training data as:
@@
Here, [bos] and [eos] are the begin-of-sequence and end-of-sequence tokens defined in the original text tokenizer. We incorporate two additional spatial tokens, [boi] and [eoi] to signify the start and end of image tokens. Furthermore, we reverse 20% of text-to-image data as captioning task to enhance the visual understanding abilities. In the scaling experiments, for each model size, we independently use 30M text-only data, 30M text-to-image data, and a mixed dataset of 60M data to train three different versions of the models. We then evaluate their performance on a series of tasks.
Since all images are tokenized into discrete image tokens that share the same vocabulary and embedding space as text tokens, the training objective is consistent with LLMs via next-token prediction and utilizes the standard cross-entropy loss. Moreover, we observe that during the continue pre-training stage, models of 7B size and above tend to encounter loss spikes in the early stages of training. It significantly impacts the training performance and convergence speed. To mitigate this issue, we reduce the max grad norm to 0.5 for larger models and use max-z loss [77] to normalize the logits, enhancing the stability of training.
Training Details. In all experiments, we initiate with a learning rate of 2e-5. Except for the scaling experiments that employed a constant learning rate, all other tests utilized a cosine learning schedule. The training batch size is set at 1024, and the max context length is set to 2048. We use the DeepSpeed Zero3 optimization strategy.
3 Scaling, Trade-offs, and Synergy in Unified Multi-modal Generation
In this section, we investigate the following questions under the unified decoder-only multi-modal paradigm of Liquid:
§3.1 Does vision generation adhere to the scaling laws observed in LLMs?
§3.2 Do vision generation and language tasks exhibit mutual interference or enhancement?
§3.3 Is there a mutually beneficial relationship between vision understanding and generation?
Figure 3: Scaling behavior of LLMs in multimodal generation across different sizes. We explore the visual generation performance of LLMs ranging from 0.5B to 32B in size after undergoing mixed training with language data and text-to-image data. The validation loss smoothly decreases as both the model size and training iterations increase, while token accuracy and VQA Score consistently rise. The VQA Score consistently increases as the validation loss decreases, indicating a strong correlation between them. Under the same training FLOPs, smaller models reach lower validation losses more quickly, but larger models ultimately achieve higher evaluation metrics.
Evaluation settings. We use validation loss, token accuracy, and VQAscore from GenAI-Bench [35] to evaluate the performance of visual generation tasks. For language ability, we use the average score across five widely-used benchmarks: HellaSwag [80], WinoGrande [56] , ARC-Easy [10], ARC-Challenge [10], and BoolQ [9]. For visual understanding tasks, we conduct evaluations on VQA-v2 [19], GQA [22], TextVQA [59], POPE [32], and MME [16].
3.1 Scaling Results on Visual Generation
We explore the visual generation performance of LLMs ranging from 0.5B to 32B in size after mixed training with 60M language and text-to-image data. As shown in Fig. 3, with the increase in model size and training iterations, the validation loss smoothly decreases, while token accuracy and VQA Score consistently increase. The VQA score continues to rise as the validation loss decreases, indicating a strong correlation between validation loss and holistic image evaluation metrics. Moreover, we observe a clear power-law scaling trend for Loss as a function of Training FLOPs, as consistent with LLMs. Under the same training FLOPs, we find smaller models achieve lower validation losses and higher VQA Scores more quickly, but larger models ultimately reach higher evaluation metrics. This may be attributed to the fact that smaller models can train for more steps under the same FLOPs, enabling them to transition more quickly from the language domain to the multimodal domain. Although smaller models can quickly acquire visual generation abilities, their upper limits are lower, making it difficult to achieve high-quality visual generation results. For further visualizations on scaling model size and training compute, please refer to Fig. 4.
Finding 1:
After equipping LLMs with visual capabilities through vocabulary expansion, their performance on visual generation tasks still follows the power-law scaling law observed in language tasks.
3.2 Is there a conflict between visual and language generation?
To explore the mutual influence between visual and language generation, we train models of six different sizes under three distinct settings and compared their performances. In addition to training on a mixed multimodal dataset of 60M mix data, we separately train two sets of models, one on 30M text-only data and the other on 30M text-to-image data, to isolate language and visual generation capabilities. We then compare the performance gaps between these isolated models and the mixed model on both language tasks and visual generation tasks.
To investigate the impact on language capabilities, we evaluate the language performance of the mixed pre-train model against the model pre-trained solely on text-only data, as illustrated in Fig. 5, we find a trade-off phenomenon when both tasks were mixed in the training of smaller models; however, as the model size increased, this trade-off gradually disappear. It confirms that larger models possess sufficient capacity to handle both visual and language spaces generation concurrently.
In language tasks, we observe a trade-off in performance of multimodal LLMs, raising the question of whether this trade-off also exists in visual generation tasks. As illustrated in Fig. 6, mixed training leads to higher validation loss on visual generation tasks for models of every size. However, the detrimental effect on the VQA-Score decreases as the model size increases. This reduction is partly due to the increased capacity of larger models and may also be attributed to the evaluation metrics approaching saturation. Nevertheless, it can be argued that as the size of the model increases, the trade-off between visual generation and language tasks gradually diminishes or even disappears. This encourages leveraging the scaling capabilities of LLMs into multimodal models.
Finding 2:
There exists a trade-off between visual and language generation, but this trade-off gradually diminishes as the model size increases.
Method Training Data Visual Generation Visual Understanding
[…]
Baseline 10M 10M 10M 19.9 0.63 0.58 0.60 60.7 51.3 39.2 75.9 909.6
Add Gen. 10M 20M 10M 12.8 0.78 0.63 0.69 64.5 53.2 39.8 78.0 1066.8
Add Und. 10M 10M 20M 14.9 0.73 0.62 0.66 63.5 53.7 40.5 76.8 1035.1
Table 1: The impact between visual understanding and generation tasks. “Visual Gen." refers to the data used for training text-guided image generation, while“Visual Und." refers to the data used for training visual understanding capabilities. Compared to the baseline, adding more understanding data improve the generation capability, and increasing the generation data also aids in understanding ability.
Figure 7: Synergistic mutual enhancement between visual understanding and generation tasks: expanding training data for one improves performance in both.
When visual inputs and outputs are modeled by LLMs in the same manner, the gap between them effectively vanishes for LLMs. In this section, we explore whether there exists a synergistic relationship between visual generation and language capabilities. We conduct three experiments. In the first set, we used a combination of 10M text-only data, 10M visual generation data, and 10M visual understanding data, resulting in a total of 30M data for the pre-training phase, which serves as the baseline. Next, we add 10M visual generation data (‘Add Gen.’) and 10M visual understanding data (‘Add Und.’) respectively. After the three pret-raining phases, we fine-tune the models on instruction data and subsequently evaluate them on both visual understanding and generation benchmarks.
As shown in Tab. 1 and Fig. 7, we observe that adding more visual understanding data during pre-training significantly improves the performance of visual generation tasks. Conversely, adding more generation data further enhances the performance of understanding tasks. This phenomenon indicates that when the modality spaces for visual understanding and generation are unified, the training of these two tasks can mutually benefit each other. Intuitively, when images are represented in the same feature space, both visual generation and visual understanding tasks require stronger consistency constraints and sufficient interaction between language and visual information. Therefore, the optimization directions for these two tasks are very similar. It further demonstrates the potential of LLMs as general-purpose multi-modal generators.
Finding 3:
There exists a synergistic relationship between visual understanding and generation, where increasing data for either leads to simultaneous improvements in both.
@@ 4 Experiments
We will obtain two types of models: the Pretrain model and the Instruction Tuning model (IT). When we initialize with an existing LLM and pretrain it on a mixed dataset of 60M multimodal data, the resulting model is referred to as the Pretrain model. This model retains fundamental language capabilities while additionally acquiring visual generation abilities. We evaluate its performance on both visual generation tasks and language tasks. For visual understanding tasks, we followed the pipeline commonly used in most visual understanding works. Specifically, we performed additional supervised fine-tuning (SFT) on the Pretrain model to enhance its instruction-following capabilities and then assessed its visual understanding performance.
@@
4.1 Quantitative Results on Visual Generation
For the image generation tasks, we evaluate the model on three benchmarks: GenAI-Bench [35], MJHQ-30K [26], and WISE [45]. GenAI-Bench [35] is an challenging image-to-text generation benchmark designed to evaluate the capabilities of visual generation models. It employs VQAScore, which leverages a visual-question-answering (VQA) model. This enables more precise evaluation of how well the generated image aligns with the text prompt, critically assessing the capability to parse scenes, objects, attributes, relationships, and engage in higher-order reasoning such as comparison and logic. MJHQ [26] calculates the Frechet Inception Distance (FID) [21] score between the generated images and 30K high-quality images to assess the quality of the generated images.
@@ 4.1 Quantitative Results on Visual Generation
Table 3: Comparison with other visual generation methods on MJHQ-30K evaluation benchmark. The FID of Liquid is lower than that of all the auto-regressive models and even outperforms most diffusion models. It indicates that the images generated by Liquid have superior aesthetic quality.
WISE [45] is the first comprehensive benchmark designed to evaluate world knowledge integration in text-to-image generation, featuring 1,000 carefully curated prompts spanning 25 diverse subdomains. It proposes WiScore as evaluation metric to systematically assesses knowledge-image alignment through a weighted combination of semantic consistency, physical realism, and aesthetic quality. These three benchmarks collectively evaluate the generated images from three perspectives: text-image alignment, image realism and fidelity, and world knowledge with reasoning.
Type Model Cultural Time Space Biology Physics Chemistry Overall
Gen. Only FLUX.1-dev 0.48 0.58 0.62 0.42 0.51 0.35 0.50
FLUX.1-schnell 0.39 0.44 0.50 0.31 0.44 0.26 0.40
PixArt-Alpha 0.45 0.50 0.48 0.49 0.56 0.34 0.47
playground-v2.5 0.49 0.58 0.55 0.43 0.48 0.33 0.49
SD-v1-5 0.34 0.35 0.32 0.28 0.29 0.21 0.32
SD-2-1 0.30 0.38 0.35 0.33 0.34 0.21 0.32
SD-XL-base-0.9 0.43 0.48 0.47 0.44 0.45 0.27 0.43
SD-3-medium 0.42 0.44 0.48 0.39 0.47 0.29 0.42
Und. & Gen. Ours 0.38 0.42 0.53 0.36 0.47 0.30 0.41
Emu3 0.34 0.45 0.48 0.41 0.45 0.27 0.39
Janus-1.3B 0.16 0.26 0.35 0.28 0.30 0.14 0.23
JanusFlow-1.3B 0.13 0.26 0.28 0.20 0.19 0.11 0.18
Janus-Pro-1B 0.20 0.28 0.45 0.24 0.32 0.16 0.26
Janus-Pro-7B 0.30 0.37 0.49 0.36 0.42 0.26 0.35
Orthus-7B-base 0.07 0.10 0.12 0.15 0.15 0.10 0.10
Orthus-7B-instruct 0.23 0.31 0.38 0.28 0.31 0.20 0.27
Show-o-demo 0.28 0.36 0.40 0.23 0.33 0.22 0.30
VILA-U-7b-256 0.26 0.33 0.37 0.35 0.39 0.23 0.31
Table 4: Normalized WiScore of different models on WISE benchmark. Liquid substantially surpasses all unified MLLMs while rivaling leading specialized visual generation models. * indicates the model underwent text-to-image fine-tuning rather than retaining full comprehension and generation capabilities. “Und.” and “Gen.” denote “understanding” and “generation”.
@@
Text-image Alignment. As shown in Tab. 4, compare with other auto-regressive based methods, Liquid achieves a better overall score under both basic prompts and advanced prompts. This suggests that the images generated by Liquid align better semantically with the input text prompts. Notably, Liquid also outperforms some well-established diffusion models like SD v2.1 [55] and SD-XL [48] for both basic and advanced prompts. Compared to these diffusion models, Liquid utilizes significantly fewer image data, which indicating that learning based on LLMs can assist the model in understanding the semantic association between the generated content and prompts, while also offering higher training efficiency. Moreover, it demonstrates that LLMs have strong potential for generating complex visual content.
@@
Image Realism and Fidelity. In Tab. 3, we report FID on MJHQ-30K to compare the images quality generated by Liquid with other models. It is observable that Liquid not only has a lower FID than all other auto-regressive methods but also surpasses most well-known diffusion models except Playground v2.5 [26], achieving a very low FID of 5.47. It indicates that LLMs are also capable of generating high-quality images, providing proof that the upper limit of LLMs in terms of image aesthetic quality is not inferior to diffusion models. Further more, due to the capability of LLM to generate dynamically-length content in the form of next-token prediction, the convenience can be applied to visual generation. We find that by appending instructions about the resolution to the input text prompt, such as "length is: width is:", the model can quickly learn to generate the corresponding code according to the specified number of rows and columns. Fig 8 demonstrates the generation results at various resolutions, showcasing the flexibility of Liquid.
World Knowledge and Reasoning. Compared to prior methods that only evaluate shallow text-image alignment, WISE introduces more challenging reasoning-driven prompts (e.g., ‘Einstein’s favorite musical instrument’), requiring models to generate images based on deeper semantic understanding. AAs shown in Tab. 4, Liquid significantly outperforms other unified MLLMs in complex reasoning scenarios for visual generation and achieves overall WiScore at 0.41, Notably, it even surpasses specialized generative models such as SD series and FLUX.1-schnell, demonstrating its strong retention of world knowledge and reasoning capabilities.
Method BaseModel Training BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c MMLU
Llama2 [70] Llama-7B 77.4 78.8 48.3 77.2 69.2 75.2 45.9 45.3
Llama-13B 81.7 80.5 50.3 80.7 72.8 77.3 49.4 54.8
Chameleon [64] Chameleon-7B 81.4 79.6 57.0 74.2 70.4 76.1 46.5 52.1
Chameleon-34B 86.0 83.3 63.3 82.7 78.5 84.1 59.7 65.8
Gemma [65] Gemma-7B 83.4 81.2 51.8 80.9 75.1 53.8 81.0 62.6
VILA-U [75] Llama2-7B multi-modal 70.6 73.5 47.7 54.2 57.3 51.6 34.0 25.5
Ours Gemma-7B text-only 82.1 81.4 46.9 77.2 71.7 76.3 48.8 55.5
Gemma-7B multi-modal 81.0 81.0 46.7 76.1 72.7 75.6 49.0 56.0
Llama3-1B text-only 61.0 75.3 41.7 62.7 59.7 61.5 34.0 27.2
Llama3-1B multi-modal 53.3 73.8 40.3 55.0 54.5 59.0 32.4 24.5
Qwen2.5-7B text-only 80.8 79.5 45.4 75.4 70.0 74.1 50.4 64.9
Qwen2.5-7B multi-modal 80.0 78.7 45.5 73.5 66.7 74.5 49.2 62.0
Qwen2.5-32B text-only 85.2 80.5 46.1 79.2 74.0 78.9 56.1 68.0
Qwen2.5-32B multi-modal 84.9 79.8 45.6 79.7 73.5 78.2 56.3 66.5